Obtain a four-term Taylor polynomial approximation valid near x = 0 for each (1+ x)-⅓. Estimate the range of x over which three term polynomials will give two decimal accuracy.


1
Expert's answer
2021-04-15T07:20:32-0400

Given, expression is-


"f(x)=(1+x)^{-\frac{1}{3}}"


On expanding the expression, we get the four term Taylor polyomial


"f(x)=1-\dfrac{1}{3}x+\dfrac{2}{9}x^2-\dfrac{14}{81}x^3"


Range i.e. Region of convergence of x is |x|<1